A hospital needs 40L of 5% solution. They have 20% and 4%. How many Liters of each 20% and 4% would they need to mix to obtain the 40L of 5% solution?
Answer by nshah11(47)

The total liters of each of the 20% and 4% solution must add up to 40.
(0.2(x) + 0.04(40-x))/40 = 0.05
0.15(x) + 1.6 = 2
x = 0.4/0.15 = 2.67 L of 20% solution.
37.33 L of 4% solution.